# Three-dimensional mapping of coronal magnetic field and plasma parameters in a solar flare

### v1

@@ 1 Introduction
A new and so far unique way of inferring evolving magnetic field directly in the flare volume is based on the microwave imaging spectroscopy of nonthermal emission produced by energetic electrons accelerated in the solar flare that produce gyrosynchrotron emission, while moving in the ambient magnetic field (Gary et al., 2013). Broadband imaging instruments such as EOVSA (Gary et al., 2018) provide spectral imaging across a broad frequency range (1–18 GHz), enabling spatially resolved measurements of magnetic field over the flaring volume (Fleishman et al., 2020).
@@ 2.4 Inferring the 3D locations of radio emission
Figure 4: Goodness of the spectral fit. Left: $\chi^{2}$ (=reduced Chi2) map for the time frame of 18:57:24 UT. Red contour shows the XRT brightness contour computed for 18:57:08 UT. Middle: Example of the spatially resolved spectrum from the pixel shown by the red cursor in the left panel and corresponding good spectral model fit shown by the blue curve. The vertical dashed red line marks the spectral peak that demarcates the optically thick (to the left) and thin (to the right) parts of the spectrum. Right: Evolving 2D histogram of the $\chi^{2}$ metric (colored dots) and the corresponding median values (black symbols with error bars). This demonstrates that the spectral fits are mainly acceptable.
@@ 3 Spectral model fitting
Here, we apply the method of microwave imaging spectroscopy (Gary et al., 2013; Fleishman et al., 2020, 2022) to infer the spatial distributions and temporal evolution of plasma parameters in the 2021-May-07 solar flare. Using EOVSA observations we constructed multifrequency images of the microwave-emitting region, forming an evolving 3D data cube in frequency and space. At each spatial pixel of the brightness maps, we fit the observed spectrum with a homogeneous gyrosynchrotron emission model using a dedicated tool called GSFIT (Fleishman et al., 2020). This uniform source represents a bar with $2\arcsec\times 2\arcsec$ front area (prescribed by the pixel size) and 8″ depth (of the order of the individual features in the flare images; see Fig. 2). For this spectral model fitting, we adopted an isotropic single power-law model for the nonthermal electrons with fixed $E_{\min}=10$ keV and $E_{\max}=5$ MeV. We also fixed the temperature of the ambient plasma at $T=20$ MK to exclude noticeable free-free absorption as our trial spectral fit runs with free temperature showed that the fit results are not sensitive to $T$ provided it is reasonably large. Therefore, the free fit parameters are: the nonthermal number density, $n_{nth}$, and power-law index, $\delta$, the thermal density, $n_{th}$, the magnetic field, $B$, and its angle, $\theta,$ to the LOS. Typically, $\theta$ is not well-constrained from the fits; however, we found that it is important to keep it free to permit enough flexibility for the fit to find a valid solution. By repeating this analysis over successive time steps, we obtained evolving parameter maps that reveal the dynamics of the flare, including the magnetic field strength, $B$, the thermal plasma density, $n_{th}$, and the parameters of the nonthermal electron distribution.
@@ 3.1 Spectral fitting results
Figure 5 illustrates instantaneous parameter maps and their evolution. The time frame, 18:57:24 UT, the same as in Fig. 4, was selected at the end of the first analyzed impulsive peak of the radio burst, when, presumably, the effect of source nonuniformity along the LOS must be minimal [^1] If we have two (or more) sources along the LOS at a decay phase, emission from one of them might decay slower than the other(s). Thus, at the late decay phase emission from that one may dominate.. All demonstrated parameter maps are rather smooth, albeit with several defects or artifacts. The thermal number density map displays a nice loop-like structure, assumed to be tracing a hot and dense post-flare loop with number density within the range of $n_{th}\sim(2-8)\times 10^{10}$ cm-3. This loop structure projects onto a bright AIA feature clearly seen in Fig. 2, which is partly occulted by a cool filament. The filament material is transparent for the microwave emission; thus, the loop structure is recovered using the microwave spectral model fitting. The nonthermal number density is considerably lower here, about $n_{nth}\sim 10^{6}$ cm-3; the spectrum of the nonthermal electrons is very soft, $\delta\gtrsim 10$, while distinctly different from a purely thermal one (see, e.g., Fleishman et al., 2015; Wang et al., 2017). Such a very tiny population of the nonthermal electrons is expected at the phase of local minimum of the microwave brightness.
@@
During the first analyzed impulsive peak, 18:54:56–18:57:24 UT, the magnetic field in this location goes down from 700 G to 470 G, with the rate of $\dot{B}\sim-7$ G s-1 similar to earlier reported values (Fleishman et al., 2020). However, this only lasts half of minute and then slowly rises up to 810 G with the rate $\dot{B}\sim 3$ G s-1. In principle, these changes in the inferred magnetic field might represent a geometrical effect, when different portions of the LOS are illuminated by the radio-emitting nonthermal electrons in different moments, rather than the evolution ”in place.” However, the inferred thermal number density remains in the range of the density inferred from the X-ray data; thus, we are likely detecting real temporal variations in $B$. Further apparent evolution of the magnetic field represents a combination of decreases and increases overall (anti)correlated with microwave brightness peaks in time (presumably, episodes of the free magnetic energy release) superimposed on a general trend of the magnetic field increase up to $\sim$900 G with the rate $\dot{B}\approx 0.2$ G s-1. During the maximum phase of the burst, $\sim$19:00–19:04 UT, the fit solutions for the magnetic field are not unique, even though the individual spectra are pretty smooth.
@@
These parameter trends discussed for a single pixel are consistent with general trends displayed by the third column showing evolving 2D histograms of the same parameters for all pixels inscribed within the red ROI. However, these histograms reveal additional features, which cannot be identified for an individual pixel. The histograms of the magnetic field and the thermal number density display a pronounced dichotomy (or even a trichotomy) of solutions. This is due to aforementioned complexity of many individual spectra that show signatures of several sources along the LOS, while the uniform source spectral model can report properties of a single source only. Given that the relative contributions of these different sources to the total spectrum can vary in space and time, the spectral fitting result fluctuate between these different sources resulting in the parameter dichotomy. We note that the effect of this dichotomy is minimal in the local minimum of the microwave brightness at 18:57:00 – 18:58:00 UT; the middle of this interval is shown by the vertical red line. This dichotomy and trichotomy are further illustrated by Fig. 6, which shows a 1D histogram of the magnetic field and thermal number density within the specified ROI during the entire analyzed flare duration. Interestingly, the nonthermal electron population does not exhibit any dichotomy and this could possibly point to at most a single acceleration source in this event.
@@ 3.2 Median parameter maps
Below, we describe how we capitalized on the fact that the thermal number density in a denser source inferred from the microwave data spans the range of $n_{\mathrm{th}}\sim(1{-}10)\times 10^{10}$ cm-3, with a median value near $5\times 10^{10}$ cm-3, which matches the density derived from the stereoscopic SXR data analysis perfectly (Ryan et al., 2024). In that case, the density varies between approximately $(5{-}7)\times 10^{10}$ cm-3 (their Fig. 6e); thus, we can focus on that denser source. This match is consistent with the assumption that the volume filling factor is close to unity in the flaring source. Given the absence of comparably dense plasma at other locations than the SXR source, we can conclude that the microwave and SXR emissions arise from the same physical volume. This provides a robust basis for assigning the 3D X-ray–derived coordinates to the plasma and magnetic field parameters inferred for our main source from the microwave analysis.
@@ 3.3 Evolution of the fit parameters moments
These data also permit us to check if the derived flare parameters are consistent with the plasma confinement by the magnetic field. At the looptop edge, we have $B=306$ G; thus, the magnetic pressure is $p_{B}=1120$ erg cm-3, which is a factor of 5 higher than the thermal pressure $p_{th}\approx 230$ erg cm-3 obtained for the microwave derived thermal number density and X-ray derived temperature of 16 MK (Ryan et al., 2024, noting that the microwave spectra are not sensitive to such high temperatures as was noted in the beginning of Sect. 3). This dominance of the magnetic pressure is sufficient for plasma confinement.
@@ 3.4 Height dependence of the magnetic field
To place our results in the context of previous observations and models, we compare them with the published profiles of the coronal magnetic field dependence on height. The green line shows an empirical dependence for a “typical” magnetic field in active regions (AR) inferred by Dulk and McLean (1978) from a combination of different indirect techniques. Most of our inferred data are located above this line. The yellow line shows a maximum magnetic field $B_{\max}(h)$ obtained from a nonlinear force-free field extrapolation devised for AR 12673 at 2017-Sep-06, when this AR displayed high photospheric (Wang et al., 2018) and record-breaking coronal magnetic field (Anfinogentov et al., 2019). By construction, the yellow line represents a (soft) upper bound of the AR magnetic field in the steady state, confirmed by a compilation of various coronal magnetic field measurements performed by Fedenev et al. (2023). [^2] An interactive version of their $B(h)$ plot can be found at https://www.leibniz-kis.de/en/research/solar-magnetism/coronal-magnetic-field/. Our inferred magnetic field data are mostly below this line.
@@
Figure 11: Maps of the Alfvén velocity (left), plasma beta (center), and plasma-to-gyro frequency ratio (right) for the high-B source (top two rows, LOS intervals 18:55:00–18:57:00 and 18:57:00–18:58:00 UT) and the low-B source (bottom row, 18:55:00–18:57:$0$ UT). The red contour shows the 20% ROI level from Hinode/XRT at 18:56:25 UT, and the violet contour indicates the 10% brightness level from the 5.79 GHz EOVSA image at 18:55:44 UT.
Figure 11 presents the spatial distributions of the Alfvén speed ($v_{\mathrm{A}}$), plasma beta ($\beta$), and the plasma-to-gyrofrequency ratio ($\alpha=f_{\mathrm{pe}}/f_{\mathrm{Be}}$) for two intervals: 18:55:00–18:57:00 UT (top row) and 18:57:00–18:58:00 UT (middle and bottom rows; the bottom row shows the results obtained for the low-$B$ source; see Sect. 3.2 and the bottom row of Fig. 7). Within the ROI (red contour), plasma beta remains well below unity in the main source, indicating a magnetically dominated environment. The Alfvén speed ranges there from $2\times 10^{8}$ to over $8\times 10^{8}\,\text{cm\,s}^{-1}$. The ratio $\alpha$ stays around 1 across most of the ROI. In contrast, in the secondary, ”low-BB” source, the Alfvén velocity is about ten times smaller than in the main source, the plasma $\beta$ is about 1, and the frequency ratio is well above 1.
@@ Acknowledgements.
This work was supported in part by NSF grant AGS-2425102, and NASA grant 80NSSC23K0090 to New Jersey Institute of Technology. TK was supported by DFG grant eBer-24-58553. SY is supported by NASA grant 80NSSC24K1242 and NSF grant AGS-2334931 to NJIT.

### v2

@@ 1 Introduction
A new and so far unique way of inferring evolving magnetic field directly in the flare volume is based on the microwave imaging spectroscopy of nonthermal emission produced by energetic electrons accelerated in the solar flare that produce gyrosynchrotron emission, while moving in the ambient magnetic field (Gary et al., 2013). Broadband imaging instruments such as EOVSA (Gary et al., 2018) provide spectral imaging across a broad frequency range (1–18 GHz), enabling spatially resolved measurements of magnetic field over the flaring volume (Fleishman et al., 2020, 2026).
@@ 2.4 Inferring the 3D locations of radio emission
Figure 4: Goodness of the spectral fit. Left: $\chi^{2}$ (=reduced Chi2) map for the time frame of 18:57:24 UT. Red contour shows the XRT brightness contour computed for 18:57:08 UT. Middle: Example of the spatially resolved spectrum from the pixel shown by the red cursor (i.e., the intersection of the two red dashed lines) in the left panel and corresponding good spectral model fit shown by the blue curve. The vertical dashed red line marks the spectral peak that demarcates the optically thick (to the left) and thin (to the right) parts of the spectrum. Right: Evolving 2D histogram of the $\chi^{2}$ metric (colored dots) and the corresponding median values (black symbols with error bars). This demonstrates that the spectral fits are mainly acceptable.
@@ 3 Spectral model fitting
Here, we apply the method of microwave imaging spectroscopy (Gary et al., 2013; Fleishman et al., 2020, 2022, 2026) to infer the spatial distributions and temporal evolution of plasma parameters in the 2021-May-07 solar flare. Using EOVSA observations we constructed multifrequency images of the microwave-emitting region, forming an evolving 3D data cube in frequency and space. At each spatial pixel of the brightness maps, we fit the observed spectrum with a homogeneous gyrosynchrotron emission model using a dedicated tool called GSFIT (Fleishman et al., 2020). This uniform source represents a bar with $2\arcsec\times 2\arcsec$ front area (prescribed by the pixel size) and 8″ depth (of the order of the individual features in the flare images; see Fig. 2). For this spectral model fitting, we adopted an isotropic single power-law model for the nonthermal electrons with fixed $E_{\min}=10$ keV and $E_{\max}=5$ MeV. We also fixed the temperature of the ambient plasma at $T=20$ MK to exclude noticeable free-free absorption as our trial spectral fit runs with free temperature showed that the fit results are not sensitive to $T$ provided it is reasonably large. Therefore, the free fit parameters are: the nonthermal number density, $n_{nth}$, and power-law index, $\delta$, the thermal density, $n_{th}$, the magnetic field, $B$, and its angle, $\theta,$ to the LOS. Typically, $\theta$ is not well-constrained from the fits; however, we found that it is important to keep it free to permit enough flexibility for the fit to find a valid solution. By repeating this analysis over successive time steps, we obtained evolving parameter maps that reveal the dynamics of the flare, including the magnetic field strength, $B$, the thermal plasma density, $n_{th}$, and the parameters of the nonthermal electron distribution.
@@ 3.1 Spectral fitting results
Figure 5 illustrates instantaneous parameter maps and their evolution. The time frame, 18:57:24 UT, the same as in Fig. 4, was selected at the end of the first analyzed impulsive peak of the radio burst, when, presumably, the effect of source nonuniformity along the LOS must be minimal [^1] If we have two (or more) sources along the LOS at a decay phase, emission from one of them might decay slower than the other(s). Thus, at the late decay phase emission from that one may dominate.. All demonstrated parameter maps are rather smooth, albeit with several defects or artifacts. The thermal number density map displays a nice loop-like structure, assumed to be tracing a hot and dense post-flare loop with number density within the range of $n_{th}\sim(2-8)\times 10^{10}$ cm-3. This loop structure projects onto a bright AIA feature clearly seen in Fig. 2, which is partly occulted by a cool filament. The filament material is transparent for the microwave emission; thus, the loop structure is recovered using the microwave spectral model fitting. The nonthermal number density in this loop is about $n_{nth}\sim 10^{6}$ cm-3; the spectrum of the nonthermal electrons is very soft, $\delta\gtrsim 10$, while distinctly different from a purely thermal one (see, e.g., Fleishman et al., 2015; Wang et al., 2017). Such a very tiny population of the nonthermal electrons is expected at the phase of local minimum of the microwave brightness.
@@
During the first analyzed impulsive peak, 18:54:56–18:57:24 UT, the magnetic field in this location goes down from 700 G to 470 G, with the rate of $\dot{B}\sim-7$ G s-1 similar to earlier reported values (Fleishman et al., 2020, 2026). However, this only lasts half of minute and then slowly rises up to 810 G with the rate $\dot{B}\sim 3$ G s-1. In principle, these changes in the inferred magnetic field might represent a geometrical effect, when different portions of the LOS are illuminated by the radio-emitting nonthermal electrons in different moments, rather than the evolution ”in place.” However, the inferred thermal number density remains in the range of the density inferred from the X-ray data; thus, we are likely detecting real temporal variations in $B$. Further apparent evolution of the magnetic field represents a combination of decreases and increases overall (anti)correlated with microwave brightness peaks in time (presumably, episodes of the free magnetic energy release) superimposed on a general trend of the magnetic field increase up to $\sim$900 G with the rate $\dot{B}\approx 0.2$ G s-1. During the maximum phase of the burst, $\sim$19:00–19:04 UT, the fit solutions for the magnetic field are not unique, even though the individual spectra are pretty smooth.
@@
These parameter trends discussed for a single pixel are consistent with general trends displayed by the third column showing evolving 2D histograms of the same parameters for all pixels inscribed within the red ROI. However, these histograms reveal additional features, which cannot be identified for an individual pixel. The histograms of the magnetic field and the thermal number density display a pronounced dichotomy (or even a trichotomy) of solutions. This is due to aforementioned complexity of many individual spectra that show signatures of several sources along the LOS, while the uniform source spectral model can report properties of a single source only. Given that the relative contributions of these different sources to the total spectrum can vary in space and time, the spectral fitting result fluctuate between these different sources yielding the parameter dichotomy. We note that the effect of this dichotomy is minimal in the local minimum of the microwave brightness at 18:57:00 – 18:58:00 UT; the middle of this interval is shown by the vertical red line. This dichotomy and trichotomy are further illustrated by Fig. 6, which shows a 1D histogram of the magnetic field and thermal number density within the specified ROI during the entire analyzed flare duration. Interestingly, the nonthermal electron population does not exhibit any dichotomy and this could possibly point to at most a single acceleration source in this event.
@@ 3.2 Median parameter maps
Below, we capitalized on the fact that the thermal number density in a denser source inferred from the microwave data spans the range of $n_{\mathrm{th}}\sim(1{-}10)\times 10^{10}$ cm-3, with a median value near $5\times 10^{10}$ cm-3, which matches the density derived from the stereoscopic SXR data analysis perfectly (Ryan et al., 2024). In that case, the density varies between approximately $(5{-}7)\times 10^{10}$ cm-3 (their Fig. 6e); thus, we can focus on that denser source. This match is consistent with the assumption that the volume filling factor is close to unity in the flaring source. Given the absence of comparably dense plasma at other locations than the SXR source, we can conclude that the microwave and SXR emissions arise from the same physical volume. This provides a robust basis for assigning the 3D X-ray–derived coordinates to the plasma and magnetic field parameters inferred for our main source from the microwave analysis.
@@ 3.3 Evolution of the fit parameters moments
These data also permit us to check if the derived flare parameters are consistent with the plasma confinement by the magnetic field. At the looptop edge, we have $B=306$ G; thus, the magnetic pressure is $p_{B}=1120$ erg cm-3, which is a factor of 5 higher than the thermal pressure $p_{th}\approx 230$ erg cm-3 obtained for the microwave derived thermal number density and X-ray derived temperature of 16 MK (Ryan et al., 2024, the microwave spectra are not sensitive to such high temperatures as was noted in the beginning of Sect. 3). This dominance of the magnetic pressure is sufficient for plasma confinement.
@@ 3.4 Height dependence of the magnetic field
To place our results in the context of previous observations and models, we compare them with the published profiles of the coronal magnetic field dependence on height. The green line shows an empirical dependence for a “typical” magnetic field in active regions (AR) inferred by Dulk and McLean (1978) from a combination of different indirect techniques. Most of our inferred data are located above this line. The yellow line shows a maximum magnetic field $B_{\max}(h)$ obtained from NLFFF extrapolation devised for AR 12673 at 2017-Sep-06, when this AR displayed high photospheric (Wang et al., 2018) and record-breaking coronal magnetic field (Anfinogentov et al., 2019). By construction, the yellow line represents a (soft) upper bound of the AR magnetic field in the steady state, confirmed by a compilation of various coronal magnetic field measurements performed by Fedenev et al. (2023). [^2] An interactive version of their $B(h)$ plot can be found at https://www.leibniz-kis.de/en/research/solar-magnetism/coronal-magnetic-field/. Our inferred magnetic field data are mostly below this line.
@@
Figure 11: Maps of the Alfvén velocity (left), plasma beta (center), and plasma-to-gyro frequency ratio (right) for the high-B source (top two rows, time intervals 18:55:00–18:57:00 and 18:57:00–18:58:00 UT) and the low-B source (bottom row, 18:55:00–18:57:$0$ UT). The red contour shows the 20% ROI level from Hinode/XRT at 18:56:25 UT, and the violet contour indicates the 10% brightness level from the 5.79 GHz EOVSA image at 18:55:44 UT.
Figure 11 presents the spatial distributions of the Alfvén speed ($v_{\mathrm{A}}$), plasma beta ($\beta$), and the plasma-to-gyrofrequency ratio ($\alpha=f_{\mathrm{pe}}/f_{\mathrm{Be}}$) for two intervals: 18:55:00–18:57:00 UT (top and bottom rows) and 18:57:00–18:58:00 UT (middle row; the bottom row shows the results obtained for the low-$B$ source; see Sect. 3.2 and the bottom row of Fig. 7). Within the ROI (red contour), plasma beta remains well below unity in the main source, indicating a magnetically dominated environment. The Alfvén speed ranges there from $2\times 10^{8}$ to over $8\times 10^{8}\,\text{cm\,s}^{-1}$. The ratio $\alpha$ stays around 1 across most of the ROI. In contrast, in the secondary, ”low-BB” source, the Alfvén velocity is about ten times smaller than in the main source, the plasma $\beta$ is about 1, and the frequency ratio is well above 1.
@@ Acknowledgements.
EOVSA was designed, built, and is now operated by the New Jersey Institute of Technology (NJIT) as a community facility. EOVSA operations are supported by NSF grant AGS-2436999 to NJIT. TK was supported by DFG grant eBer-24-58553. SY is supported by NASA grant 80NSSC24K1242 and NSF grant AGS-2334931 to NJIT. GF was supported in part by NSF grant AGS-2425102, and NASA grant 80NSSC23K0090 to New Jersey Institute of Technology.